OpenStudy (anonymous):

A basketball team played a total of 13 games and lost "L" games. They won five times as many games as they lost. One game was a tie. How many games did the team lose? 13/(6L) 13 13+1/(5L) 2 5L

OpenStudy (anonymous):

I think it is B

jimthompson5910 (jim_thompson5910):

Let W = # of games won So because "They won five times as many games as they lost.", we know that W = 5L Since one game was a tie, we know that they played W+L + 1 games total. But this total is 13, so W+L+1 = 13 Now plug in W = 5L to get 5L+L+1 = 13 and then solve for L 5L+L+1 = 13 6L+1 = 13 6L = 12 L = 2 So they lost 2 games (and won 2*5=10 games).

OpenStudy (anonymous):

Ok! Which option would that be though??

jimthompson5910 (jim_thompson5910):

No, how can it be 13 games lost? They played 13 games total and tied one game...so they didn't lose all 13 games.

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Thats what i meant..lol

jimthompson5910 (jim_thompson5910):

You got it